Almarai has just set the company dividend policy at 0.45 riyal per year. The company plans to be in business forever.

What is the price of this stock if the investor wants a return of 5%?

Nadec has a current annual cash dividend policy of 4.00 riyal. The price of the stock is set to

yield a return of 8%. What is the price of this stock if the dividend will be paid for 10

years and then the company repurchase the stock for 25 riyal?

SABIC has just paid its shareholders an annual dividend of 0.75 riyals and announced that the dividends would grow at an annual rate of 4% forever. If investors expect to earn an annual rate of return of 7% on this investment, how much would SABIC stock will be worth in 5 years?
Answer #1

Price of this Stock if the investor wants 9% return =Dividend/Required Rate =0.45/5% =9

Dividend =4
Number of years of dividend =10
Rate =8%
Price in year 10 =25
Price of Stock today =PMT*((1-(1+r)^-n)/r)+Price in year 10/(1+r)^n =4*((1-(1+8%)^-10)/8%)+25/(1+8%)^10=38.42

Dividend year 0 =0.75
Value of Stock in Year 5=Dividend Year 0*(1+growth)^6/(Required Rate-Growth) =0.75*(1+4%)^6/(7%-4%)=31.63